hello,
how do i simplify
2 radical 2 + 3 radical 8

Answer by rothauserc(4718)

2 radical 2 + 3 radical 8 =
2 radical 2 + 6 radical 2 =
2 radical 2 (1 + 3) =
8 radical 2